# Embedded coherent structures from MHD to sub-ion scales in turbulent solar wind at 0.17 AU

### v2

@@ 5.2 Current sheets
Current sheets are planar coherent structures that separate the plasma with different magnetic field directions. MHD classification of current sheets include rotational (RDs) and tangential (TDs) discontinuities (e.g., Baumjohann & Treumann, 1997; Tsurutani et al., 2011).
@@ 5.2.2 Tangential discontinuity
Figure 6 column (d) shows the crossing of the tangential discontinuity. The details of the model are described in appendix A.2.2. Independently on the crossing trajectory the polarisation is linear, $\lambda_{2}/\lambda_{1}\sim 0$, $\lambda_{3}/\lambda_{2}\sim 1$. Only the maximum MVA eigenvector ($\mathbf{e}_{1}$) is unambiguously defined, it is tangential to the discontinuity plane ($\mathbf{e}_{1}=\mathbf{y}\perp\mathbf{x}$). The intermediate ($\mathbf{e}_{2}$) and minimum ($\mathbf{e}_{3}$) eigenvectors are in $x-z$ plane, where $\mathbf{x}$ and $\mathbf{z}$ are normal and guide-field directions correspondingly. In this tangential discontinuity model $B_{x}=0$ $B_{z}=const$, so $\delta B_{x}=\delta B_{z}=0$. Consequently, MVA analysis can’t distinguish between the normal ($\mathbf{x}$) and the guide field ($\mathbf{z}$) directions. In general, the tangential discontinuity can be asymmetric, separating plasmas with different |B|. Thus, the tangential discontinuity can be a compressible structure. But we are using the simple model, where the value of the magnetic field modulus |B| is the same on both sides of the discontinuity.
@@ 5.3 Magnetic holes
Magnetic holes are compressible coherent structures, characterised by a localised decrease of the magnetic field modulus. Figure 6, column (e) shows the crossing of the magnetic hole model. The details of the model and some observational properties of magnetic holes are discussed in the appendix A.3. For any crossing trajectory, the polarisation is linear, $\lambda_{2}/\lambda_{1}\sim 0$, $\lambda_{3}/\lambda_{2}\sim 1$.
@@ 6 Examples of the observed structures
We consider coherent structures detected by the integrated LIM over all scales and above the threshold, $I>I_{threshold}$ (see Section 4.1). Among nearly $\sim 10^{4}$ events we have selected 374 with $I/I_{threshold}\geq 6$ for visual examination. All of them have a localised event at sub-ion scales, which is embedded in a larger event at ion scales. At its turn, this ion scale event is embedded in an MHD scale event. This successive embedding, that is shown here for the first time, is like the organisation of Russian dolls.
@@
Figure 7: Four examples of events detected on 6 November 2018 with integrated LIM $I/I_{threshold}>6$ are shown in columns. The central time of each structure $t_{0}$ is indicated in the title. Top row: Raw magnetic field in RTN reference frame. Rows 2 through 4: the bandpass filtered magnetic fluctuations at MHD, ion and sub-ion frequency ranges in local MVA reference frames. At MHD scales these structures represent: a Tangential discontinuity (example 1), an Alfvén vortex (example 2), a rotational discontinuity at the switchback boundary (example 3), and two neighboring switchbacks (example 4). In appendix B the additional information (magnetic fluctuation polarisation, alfvenicity, plasma parameters) is provided for each event. This information complement the interpretation of the coherent structures. At ion and sub-ion scales magnetic fluctuations can be interpreted as vortex like structures for the 4 examples.
@@
In Figure 7 we show 4 such examples of different types of the structures at different scales, found among the subset of 374 events. Here we restrict ourselves to show the magnetic fluctuations only. But the complementary information for each event is presented in appendix B. It shows the variation of the plasma parameters ($N_{e},T_{e},T_{p},|V|$) across the structure, and the alfvénicity.
@@
We first consider the event in the left column of Figure 7 (a-d). Panel (a) shows the low-pass filtered magnetic field in RTN reference frame during ±$\pm 500$ seconds around the central time with $t_{0}=[00:36:27]$ UT (on 6 November 2018) subtracted to the observed time. It is intended to show the global environment around the event. Using the mean observed velocity, it corresponds to a scale $\sim 3\times 10^{5}$ km, so of the order of injection scales (Chen et al., 2020). In this case the coherent structure, around $t=0$ is associated to a discontinuity but it is not associated to a switchback boundary, since $B_{R}$ is not reversing.
@@
In Figure 7 (b-d) we show the filtered magnetic field data $\delta B_{j}$ at time scales defined by Equation (7), with $j=$‘MHD’, ‘ion’ and ‘subion’. We use local MVA reference frame adapted to each scale range shown (Sonnerup & Scheible, 1998). The basis vectors $(\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3})$ are directed along the maximum, intermediate and minimum variance of the magnetic field.
@@
In Figure 7 (b), a high amplitude current sheet is observed at MHD scales. In appendix B.1 we use the plasma data to provide a physical analysis as summarized below. In Figure 11 (i-k) we show that the Walen relation $\Delta\mathbf{V}=\pm\Delta\mathbf{B}/\sqrt{4\pi\rho}$ for rotational discontinuities is violated in this example, and the pressure anisotropy correction (Hudson, 1970) is insufficient to explain the observation. In addition, the magnetic field magnitude is changing across the sheet, see the black curve in Figure 11 (l). The normalised amplitude of the jump is $\Delta B/B_{0}=0.1$, where $\Delta B=|B(t_{0}+\Delta t^{\prime})|-|B(t_{0}-\Delta t^{\prime})|=10$ nT, with $\Delta t^{\prime}=3$ s, and $B_{0}=\langle|B|\rangle_{t-t_{0}\in(-100,100)\text{ s}}=100$ nT. This localised jump is $\approx 2$ times greater than the standard deviation of $\Delta B/B_{0}$ in the 5 h interval. This is also incompatible with the rotational discontinuity, across which the magnetic field magnitude is constant. In contrast, the tangential discontinuity permits a change of the magnetic field magnitude. So we interpret this structure as a tangential discontinuity. The thickness of this current sheet can be estimated as $d=V\Delta t=480$ km, or $100$ $\rho_{i}$ and $40$ $d_{i}$, where $\Delta t=1.5$ s is the duration of the current sheet crossing, local $V=340$ km/s.
@@
In Figure 7 (e-h) the second event around $t_{0}=[01:19:20]$ UT is shown in the same format as the first event. Radial magnetic field component $B_{R}$ is negative in the center of the structure and during the time interval ±$\pm 100$ s (MHD scales). So the MHD-scale coherent structure, shown in Figure 7(f), is not associated to a switchback. Its time profile is consistent with the monopole Alfvén vortex crossing close to its center, see the black trajectory in Figure 6 (a). The velocity and magnetic field fluctuations are correlated with a proportionality coefficient $\xi$, $\delta B/B_{0}=\xi\delta V/V_{A}$, where $\xi=0.86<1$, see appendix B.2, Figure 12 (i-k). In the model of Alfvén vortex, $\xi$ is a free parameter, simply related to other parameters of the vortex: $\xi=u/\alpha$, where $\alpha$ is the inclination angle with respect to the background magnetic field, and $u=V_{vortex}/v_{th,\perp}$ is the normalised vortex propagation speed in the perpendicular plane, $v_{th,\perp}=\sqrt{2k_{B}T_{p,\perp}/m_{p}}$ is the perpendicular proton thermal speed.
@@
The direction of the Alfven vortex axis can be estimated using MVA if the vortex is crossed by the satellite off-centre. In this case, the vortex axis is directed along the direction of the minimum MVA direction $\mathbf{e}_{3}$. In the considered example the Alfvén vortex is crossed near the centre, the direction of $\mathbf{e}_{3}$ (as well as $\mathbf{e}_{2}$) is not reliably determined. However, if we assume that $\mathbf{e}_{3}$ approximate well the direction of the vortex axis, the inclination angle is $\alpha=4^{\circ}$. Then one can estimate the vortex propagation velocity $u=\xi\cdot\alpha=0.86\cdot(4/180\cdot\pi)=0.06$. So with proton thermal speed $v_{th,\perp}\simeq 50$ km/s the propagation velocity of the vortex $V_{vortex}=3$ km/s. This is negligible in comparison with the local bulk solar wind speed $V=330$ km/s, so the vortex is mainly convected across the satellite. Thus, we use the Taylor hypothesis to estimate its cross-section scale $d=V\Delta t\sin\Theta_{BV}=2.4\cdot 10^{4}$ km, or $5\cdot 10^{3}$ $\rho_{i}$ and $2\cdot 10^{3}$ $d_{i}$, where $\Delta t=80$ s is the duration of the structure, local $\Theta_{BV}=70^{\circ}$.
@@
Within this monopole Alfvén vortex, we observe smaller scale embedded vortices, see Figure 7 (c-d). The schematic sketch illustrating this embedding is presented in Figure 10 (b). Note, that the sketch is not to scale, the diameters of the embedded vortices differ by a factor of 10: $d=440$ km $=40d_{i}$ $=90\rho_{i}$ and $d=47$ km $=4d_{i}$ $=9\rho_{i}$.
@@
Figure 7 (i-l) shows a third event around $t_{0}=[02:19:38]$ UT. In the panel (i) $B_{R}$ changes sign across the sheet, meaning that the sheet forms the boundary of a switchback (see the complementary data in Figure 13 , and Appendix B.3) similarly to observations by Krasnoselskikh et al. (2020). In the panel 7 (j) the current sheet is located at $t-t_{0}=20$ s. The velocity and magnetic field jumps satisfy the Walen relation and the magnetic field magnitude is constant within the short interval $t-t_{0}\in(0,40)$ s near the center of the sheet. So we conclude that this discontinuity is rotational.
@@
Ion scale magnetic fluctuations are shown in the Figure 7 (k). The minimum variance direction $\mathbf{e_{3}}$ is along the local mean magnetic field $\mathbf{B}_{0}=\langle\mathbf{B}\rangle_{t-t_{0}\in(-0.5,0.5)\text{ s}}$. The maximum ($\delta B_{1}$) and intermediate ($\delta B_{2}$) MVA components are perpendicular to $B_{0}$ and have similar amplitudes. At the center of the structure ($t-t_{0}=0$) $\delta B_{1}$ changes sign, and $\delta B_{2}$ has a peak. So, magnetic field fluctuations are in accordance with the off-center crossing of the monopole Alfvén vortex, see blue trajectory in Figure 6(a). The cross-section scale of the vortex is $150$ km, or $25~{}\rho_{i}$ and $12~{}d_{i}$.
@@
Sub-ion scale structure (Figure 7 (l)) has typical properties for structures at sub-ion scales in our statistics: $\delta B_{1}$ has a Mexican hat-like shape, and this event has a significant compressibility $\delta|B|\sim 0.5\,\delta B_{1}$. It is very similar to the sub-ion scale structure shown in Figure 7 (d). The cross-section scale of this structure is 24 km, or 4 $\rho_{i}$ and 2 $d_{i}$. Such localized compressible magnetic fluctuations at ion scales can be interpreted as the ion Alfvén vortex of Jovanović et al. (2020).
@@
Figure 7 (m-p) shows the fourth example around $t_{0}=[02:54:08]$ UT. Our detection method catch two neighboring switchbacks, lasting 200 seconds at MHD scales. Indeed, we see in the panel (m) that $B_{R}>0$ during two time intervals in the center, and the rest of the time $B_{R}<0$. Figure 7(n) show four current sheets at the boundaries of these switchbacks (panel (n)). The central times of the four current sheets are $t-t_{0}=\{-29,0,25,85\}$ s. The complementary information is presented in appendix B.4.
@@
At ion scales (Figure 7(o)) we observe an embedded coherent structure that might represent a monopole Alfvén vortex crossed through the center. The cross-section scale is $d=160$ km $=13$ $d_{i}=23~{}\rho_{i}$. We observe a strong peak-like fluctuation of $\delta B_{1}$ at sub-ion scales (see panel (p)). The intermediate MVA fluctuation $\delta B_{2}$ is localised in the center of the event. The profile of $\delta B_{2}$ is closely similar to the one of $\delta|B|$. So this structure is compressible $\delta|B|\sim 0.2\,\delta B_{1}$. The cross-section scale is $d=12$ km $\simeq d_{i}=1.7\rho_{i}$. It belongs to the same typical class of sub-ion scale coherent structures (compressible ion scales Alfvén vortices) as in other examples.

### v4

@@ 5.2 Current sheets
Current sheets are planar coherent structures that separate the plasma with different magnetic field directions. MHD classification of current sheets includes rotational (RDs) and tangential (TDs) discontinuities (e.g., Baumjohann & Treumann, 1997; Tsurutani et al., 2011).
@@ 5.2.2 Tangential discontinuity
Figure 6 column (d) shows the crossing of the tangential discontinuity. The details of the model are described in appendix A.2.2. Independently on the crossing trajectory the polarisation is linear, $\lambda_{2}/\lambda_{1}\sim 0$, $\lambda_{3}/\lambda_{2}\sim 1$. Only the maximum MVA eigenvector $\mathbf{e}_{1}$ is unambiguously defined, it is tangential to the discontinuity plane ($\mathbf{e}_{1}=\mathbf{y}$). The intermediate ($\mathbf{e}_{2}$) and minimum ($\mathbf{e}_{3}$) eigenvectors are in $x-z$ plane, where $\mathbf{x}$ and $\mathbf{z}$ are normal and guide-field directions correspondingly. In this tangential discontinuity model $B_{x}=0$ $B_{z}=const$, so $\delta B_{x}=\delta B_{z}=0$. Consequently, MVA analysis can’t distinguish between the normal ($\mathbf{x}$) and the guide field ($\mathbf{z}$) directions. In general, the tangential discontinuity can be asymmetric, separating plasmas with different $|\mathbf{B}|$. Thus, the tangential discontinuity can be a compressible structure. Here, we use the simple model, where the value of the magnetic field modulus $|\mathbf{B}|$ is the same on both sides of the discontinuity.
@@ 5.3 Magnetic holes
Magnetic holes are compressible coherent structures, characterized by a localized decrease of the magnetic field modulus. Figure 6, column (e) shows the crossing of the magnetic hole model. The details of the model and some observational properties of magnetic holes are discussed in the appendix A.3. For any crossing trajectory, the polarisation is linear, $\lambda_{2}/\lambda_{1}\sim 0$, $\lambda_{3}/\lambda_{2}\sim 1$.
@@ 6 Examples of the observed structures
We consider coherent structures detected by the integrated LIM over all scales and above the threshold, $I>I_{threshold}$ (see Section 4.1). Among nearly $\sim 10^{4}$ events we have selected 374 with $I/I_{threshold}\geq 6$ for visual examination. All of them have a localized event at sub-ion scales, which is embedded in a larger event at ion scales. In its turn, this ion scale event is embedded in an MHD scale event. This successive embedding, which is shown here for the first time, is like the organization of Russian dolls.
@@
Figure 7: Four examples of events detected on 6 November 2018 with integrated LIM $I/I_{threshold}>6$ are shown in columns. The central time of each structure $t_{0}$ is indicated in the title. Top row: Raw magnetic field in the RTN reference frame. Rows 2 through 4: the bandpass filtered magnetic fluctuations at MHD, ion, and sub-ion frequency ranges in local MVA reference frames. At MHD scales these structures represent: a Tangential discontinuity (example 1), an Alfvén vortex (example 2), a rotational discontinuity at the switchback boundary (example 3), and two neighboring switchbacks (example 4). In appendix B the additional information (magnetic fluctuation polarisation, alfvenicity, plasma parameters) is provided for each event. This information complements the interpretation of the coherent structures. At ion and sub-ion scales magnetic fluctuations can be interpreted as vortex-like structures for the 4 examples.
@@
In Figure 7 we show 4 such examples of different types of structures at different scales, found among the subset of 374 events. Here we restrict ourselves to show the magnetic fluctuations only. But the complementary information for each event is presented in appendix B. It shows the variation of the plasma parameters ($N_{e},T_{e},T_{p},|V|$) across the structure and the alfvénicity.
@@
We first consider the event in the left column of Figure 7 (a-d). Panel (a) shows the low-pass filtered magnetic field in the RTN reference frame during ±$\pm 100$ s around the central time with $t_{0}=[00:36:27]$ UT (on 6 November 2018). Using the mean observed velocity $\sim 340$ km/s, time-scale of $200$ s corresponds to a space-scale $\sim 7\cdot 10^{4}$ km. In this case, the coherent structure, around $(t-t_{0})=0$ is associated with a discontinuity but it is not associated with a switchback boundary, since $B_{R}$ is not reversing.
@@
In Figure 7 (b-d) we show the filtered magnetic field data $\delta B_{j}$ at time scales defined by Equation (7), with $j=$‘MHD’, ‘ion’ and ‘subion’. We use the local MVA reference frame (Sonnerup & Scheible, 1998) adapted to each scale range shown. The basis vectors $(\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3})$ are directed along the maximum, intermediate, and minimum variance of the magnetic field.
@@
In Figure 7 (b), a high amplitude current sheet is observed at MHD scales. In appendix B.1 we use the plasma data to provide a physical analysis as summarized below. In Figure 11 (i-k) we show that the Walen relation $\Delta\mathbf{V}=\pm\Delta\mathbf{B}/\sqrt{4\pi\rho}$ for rotational discontinuities is violated in this example, and the pressure anisotropy correction (Hudson, 1970) is insufficient to explain the observation. In addition, the magnetic field magnitude is changing across the sheet, see the black curve in Figure 11 (l). The normalized amplitude of the jump is $\Delta B/B_{0}=0.1$, where $\Delta B=|B(t_{0}+\Delta t^{\prime})|-|B(t_{0}-\Delta t^{\prime})|=10$ nT, with $\Delta t^{\prime}=3$ s, and $B_{0}=\langle|B|\rangle_{t-t_{0}\in(-100,100)\text{ s}}=100$ nT. This localized jump is $\approx 2$ times greater than the standard deviation of $\Delta B/B_{0}$ in the 5 h interval. This is also incompatible with the rotational discontinuity, across which the magnetic field magnitude is constant. In contrast, a change in the magnitude of the magnetic field is possible when crossing a tangential discontinuity. So we interpret this structure as a tangential discontinuity. The thickness of this current sheet can be estimated as $d=V\Delta t=480$ km, or $100$ $\rho_{i}$ and $40$ $d_{i}$, where $\Delta t=1.5$ s is the duration of the current sheet crossing, local $V=340$ km/s.
@@
In Figure 7 (e-h) the second event around $t_{0}=[01:19:20]$ UT is shown in the same format as the first event. Radial magnetic field component $B_{R}$ is negative in the center of the structure and during the time interval ±$\pm 100$ s (MHD scales). So the MHD-scale coherent structure, shown in Figure 7(f), is not associated with a switchback. Its time profile is consistent with the monopole Alfvén vortex crossing close to its center, see the black trajectory in Figure 6 (a). The velocity and magnetic field fluctuations are correlated with a proportionality coefficient $\xi$, $\delta B/B_{0}=\xi\delta V/V_{A}$, where $\xi=0.86<1$, see appendix B.2, Figure 12 (i-k). In the model of Alfvén vortex, $\xi$ is a free parameter, simply related to other parameters of the vortex: $\xi=u/\alpha$, where $\alpha$ is the inclination angle of the vortex axis with respect to the background magnetic field, and $u=V_{vortex}/V_{th,\perp}$ is the normalized vortex propagation speed in the vortex plane.
@@
The direction of the Alfvén vortex axis can be estimated using MVA if the vortex is crossed by the satellite off-center. In this case, the vortex axis is directed along the direction of the minimum MVA direction $\mathbf{e}_{3}$. In the considered example the Alfvén vortex is crossed near the center, the direction of $\mathbf{e}_{3}$ (as well as $\mathbf{e}_{2}$) is not reliably determined. However, if we assume that $\mathbf{e}_{3}$ approximates well the direction of the vortex axis, the inclination angle is $\alpha=4^{\circ}$. Then one can estimate the vortex propagation velocity $u=\xi\cdot\alpha=0.86\cdot(4/180\cdot\pi)=0.06$. Thus, with a local proton thermal speed of $v_{th,\perp}\simeq 50$ km/s, the vortex propagation velocity is $V_{vortex}=3$ km/s. This is negligible in comparison with the local bulk solar wind speed $V=330$ km/s, so the vortex is mainly convected across the satellite. Thus, we can use the Taylor hypothesis to estimate its cross-section scale: $d=V\Delta t\sin\Theta_{BV}=2.4\cdot 10^{4}$ km, or $5\cdot 10^{3}\rho_{i}$ and $2\cdot 10^{3}d_{i}$, where $\Delta t=80$ s is the duration of the structure and $\Theta_{BV}=70^{\circ}$ is the local field-to-flow angle.
@@
Within this large-scale monopole Alfvén vortex, we observe smaller embedded vortices, see Figure 7(g,h). The schematic sketch illustrating this embedding is presented in Figure 10(b). Note, that the sketch is not to scale, the diameters of the embedded vortices differ by a factor of 10: $d=440$ km $=90\rho_{i}=40d_{i}$ at scales that we call ion scales and $d=47$ km $=9\rho_{i}=4d_{i}$, at so-called sub-ion scales.
@@
Figure 7(i-l) shows a third event around $t_{0}=[02:19:38]$ UT. In the panel (i), $B_{R}$ changes sign across the sheet, meaning that the sheet forms the boundary of a switchback (see the complementary data in Figure 13 , and Appendix B.3) similarly to observations by Krasnoselskikh et al. (2020). The current sheet is located at $t-t_{0}=20$ s, see Figure 7(j). The velocity and magnetic field jumps satisfy the Walen relation and the magnetic field magnitude is constant within the short interval $t-t_{0}\in(0,40)$ s near the center of the sheet. So we conclude that this discontinuity is rotational. The thickness of the current sheet ($\sim 40$ s), in terms of spatial scales, corresponds to $1.4\cdot 10^{4}$ km $=2\cdot 10^{3}\rho_{i}=1.2\cdot 10^{3}d_{i}$.
@@
Ion scale magnetic fluctuations are shown in Figure 7(k). The minimum variance direction $\mathbf{e_{3}}$ is along the local mean magnetic field $\mathbf{B}_{0}=\langle\mathbf{B}\rangle_{t-t_{0}\in(-0.5,0.5)\text{ s}}$. The maximum ($\delta B_{1}$) and intermediate ($\delta B_{2}$) MVA components are perpendicular to $B_{0}$ and have similar amplitudes. At the center of the structure ($t-t_{0}=0$) $\delta B_{1}$ changes sign, and $\delta B_{2}$ has a peak. So, magnetic field fluctuations are in accordance with the off-center crossing of the monopole Alfvén vortex, see the blue trajectory in Figure 6(a). The cross-section scale of the vortex is $150$ km, or $\sim 25\rho_{i}$ and $12d_{i}$.
@@
Sub-ion scale structure, Figure 7 (l), has typical properties for structures at sub-ion scales in our statistics: $\delta B_{1}$ has a Mexican hat-like shape, and this event has a significant compressibility $\delta|B|\sim 0.5\,\delta B_{1}$. It is very similar to the sub-ion scale structure shown in Figure 7 (d). The cross-section scale of this structure is 24 km, or 4 $\rho_{i}$ and 2 $d_{i}$. Such localized compressible magnetic fluctuations at ion scales can be interpreted as the ion Alfvén vortex of Jovanović et al. (2020).
@@
Figure 7 (m-p) shows the fourth example around $t_{0}=[02:54:08]$ UT. Our detection method catches two neighboring switchbacks, lasting 200 seconds at MHD scales. Indeed, we see in panel (m) that $B_{R}>0$ during two time intervals in the center, and the rest of the time $B_{R}<0$. Figure 7(n) shows four current sheets at the boundaries of these switchbacks. The central times of the four current sheets are $t-t_{0}=\{-29,0,25,85\}$ s. The complementary information is presented in appendix B.4.
@@
At ion scales, Figure 7(o), we observe an embedded coherent structure that might represent a monopole Alfvén vortex crossed through the center. The cross-section scale is $d=160$ km $=23\rho_{i}=13d_{i}$. We observe a strong peak-like fluctuation of $\delta B_{1}$ at sub-ion scales (see panel (p)). The intermediate MVA fluctuation $\delta B_{2}$ is localized in the center of the event. The profile of $\delta B_{2}$ is closely similar to the one of $\delta|B|$. So this structure is compressible $\delta|B|\sim 0.2\,\delta B_{1}$. The cross-section scale is $d=12$ km $\simeq 1.7\rho_{i}=d_{i}$. It belongs to the same typical class of sub-ion scale coherent structures (compressible ion scales Alfvén vortices) as in other examples.